EE 290 G  Intro to MEMS  Spring 1997
Instructor  Kristofer S.J. Pister

pister@eecs ← easiest method of contact

512 Cory

Office hours  M 11-12, W 2-3, and by appointment (ideally email)
Required texts:  none

selected papers will be distributed

Suggested references (in no particular order, and of questionable relevance):

Transducers 91 ($15 if you want one)

Gray and Meyer, Analog Integrated Circuits

Ristic, Sensor Technology and Devices

Weste and Eshraghian, Principles of CMOS VLSI design

Grading

Homework 30%
Midterm  15% probably Thursday, 3/20/95, in class
Project  30%
Final exam 25%  Tuesday, 3/19/95, 11:30 am
Homework  Will emphasize design, layout, and simulation of devices and systems.

Collaboration is encouraged.

Project  Writeup is due last week of class. Four pages max, two column format,

no exceptions.

Oral presentations will be required.

Collaboration is encouraged.

Homework 1 - Due Monday before midnight!!

Send me email from any Unix account with the following info:

• Name

• student ID

• department

• layout, design, or CAD software you have used (if any). Separate multiple program names with commas (if necessary).

• CAD software you have access to.

Place each answer on a separate line. For example:

Oski T. Student
123456789
M.E.
L-edit, MacLayout
none

Note: This is a real homework assignment, intended in part to see if you can follow simple directions, and to get you used to submitting homework electronically. Based on past experience there will be some of you who ignore the simple instructions above. Sigh.